Other formats

    Adobe Portable Document Format file (facsimile images)   TEI XML file   ePub eBook file  


    mail icontwitter iconBlogspot iconrss icon

Victoria University Antarctic Research Expedition Science and Logistics Reports 1984-85: VUWAE 29


page 37


AUGUST Saturday 25 - Sunday 26

Reconnaissance to Butter Point (Saxby, Pyne, Macpherson).

Route: McMurdo Ice Shelf to tip of the shelf; cracks in sea ice off ice shelf did not require bridging for Snotrac 34. Cracks had not spread significantly since Saxby's reconnaissance of 18 August and had refrozen further. Widest crack approximately 1.5 m was crossed about 2 km west of ice shelf tip.

Good smooth ice to Strand Moraines except 3 km wide band of moderately rough ice near the moraines that stretches north past Butter Point. Smooth ice corridor inland of rough ice - good travel to Butter Point.

Butter Point Camp: Heavy snow, drifts as high as buildings on lee side (north). Occupied mess on Saturday night, checked most buildings before leaving Sunday.

Route: Similar route to before except tried to cut through rough ice band nearer north end of Strand Moraines. Not much improvement. Largest ice crack at ice shelf required bridging and had widened about 1 m since previous day. Return on ice shelf route.

Sunday 26

Microscopes: Checked and tested Scott Base stereo and polarising microscopes for use in the drill site science lab.

Tuesday 28

Sea ice route - Scott Base to McMurdo Ice Shelf Tip (Pyne, Macpherson, Stephenson, Brown, Ross).

Good ice route, small low broken ice patches (less than 0.5 m high) closest to McMurdo. Sea ice thickness to ice shelf varied from 32" - 33". Several cracks were mapped around the ice shelf tip. The large crack crossed on Butter Point recee (25-26 August) had grown to a constant width of 4-5 m. Return to Scott Base flagging a possible route. Thickness of ice at edge of large crack was 30".

Survival course for Barrett, Ashby, Ward, Wright, run by Saxby.

Thursday 30

Survival course for all other CIROS personnel run by Macpherson and Fry (Scott Base technician). (Postponed from Wednesday). Overnight, included shelter building, tent erection and elementary snowcraft.

Friday 31

Sea ice route and crack system at Ice Shelf Tip (Barrett, Gupwell, Vardy, Saxby).

Check of crack system showed that the largest crack had closed at the dogleg with two possible crossing sites. The smaller cracks near the ice shelf, however, would be more difficult for the D4. The crack off McMurdo could be bridged with snow.

Sunday September 2

Sea ice route (Barrett, Griffith, Topping, Macpherson, Stephenson).

Cracks at the ice shelf were continuing to freeze up and close. The ice shelf route can be used so that only the McMurdo crack and Barrett's crack need be crossed. The D4 therefore should be able to go to Butter Point.

August 27 - September 2

Science personnel at Scott Base unpacked science cargo and began checking, testing and modification of equipment for the drill site science lab.

page 38

All science cargo arrived on the Winfly flights with the exception of the Helium gas and regulator ordered by Antarctic Division.

Surveying equipment checked and calculations begun for drill site surveying.

Analysis of MSSTS core logs begun by Barrett for comparison with CIROS, etc.

Science personnel assisted with other preparation for CIROS at Scott Base.

Monday 3

Preparation of equipment for advance party to Butter Point.

Tuesday 4

Advance party (Pyne, Stephenson, Neale, Miller, Hoburn) left for Butter Point with Snotrac 34 and toboggans 39, 45. Fuel problems with 39 which was left at the ice shelf to be returned to Scott Base.

Poor visibility and weather, partly cleared near Strand Moraines. Dropped track (Snotrac 34) about 1600 and decided to camp 3 km from moraines.

D4 left Scott with cargo, escort Dodge. Left D4 and cargo on ice shelf returned Dodge and personnel to Scott for night.

Wednesday 5

Advance party continue to Butter Point arrive at 1430. Generator running by 1630. D4 arrive at Butter Point about 2000 hrs escort by Dodge. D4 and Dodge remain on sea ice for night. Access onto Butter Point requires work with snow bridge over tide crack.

Friday 7

Building access ramp onto Batter Point. Large sea ice block at access on glacier required blasting. Slab 3 m × 2 m × 1-1.5 m thick.

Ice Quarry south of Camp blasted by Pyne and Wright. Snow 0.5-0.7 m deep cleared of ice and trench blasted for D4 to clear and expose a working face.

Ice strongly shattered after blast, easily removed with D4. Future blasting for ice blocks try 5′ holes with 1/2 stick at bottom, single row, holes about 1 m apart. For future quarries the trench should be aligned parallel to the SSE wind direction for easier clearing of snow drifts by the D4.

page 39

Saturday 8

Dug out Science Lab fox inspection at Butter Point.

Sunday 9

Drill site survey (Pyne, Neale) using Hilux 48 in poor weather conditions. Surveyed 8 km from drill site before trigs became clouded. Trig Herb beacon appears to have blown over.

September 3 - September 9

Science personnel at Scott Base continued working on science equipment.

Personnel at Butter Point assisted with camp facilities.

Monday 10

Poor weather, but attempted to continue drill site location with Dodge, however, bad fuel (Mogas) at Butter Point meant did not get away from camp.

Friday 14

Attempt to climb Trig Herb to re-erect survey beacon for drill site surveying. Snotrac 34 (Macpherson, Pilcher, McLennan) could not get to the foot of Herb on the Bowers Piedmont Glacier. Stuck in small crevasse, rescued by Pyne, Nealeon IH500 tractor.

Thursday 13

Located drill site approximately 500 m west of SP 130 position. Large flat ice area - ice thickness 3′11″. Precise location needs Trig Herb beacon.

Saturday 15

Reset old USGS survey mark on Herb. Drum beacon blown off. (Macpherson, Neale, Pilcher, McLennan). Climbed up Herb in 3 hours from sea ice.

Sunday 16

Resurvey drill site area. (Pyne, Neale, Wright). Area found on 13th looks to be the best position of a large flat ice area.

Attempt to survey ice edge not completed because of rough ice further east. Ice edge estimated 5-8 km east of drill site area. Breakout could be easily seen from Herb and is seaward (1-2 km) of DVDP 15 berg and is estimated to be 12 km east of Bowers Piedmont (half way between Butter Point and Strand Moraines).


Science personnel assisted around camp. Geology trip to Explorers Cove in New Harbour lead by Barrett on 16th.

Monday 17

Erected prism set on C. Bernacchi for sea ice movement (Neale, Wright). Pyne, Macpherson continued to Marble Point in Hilux 48. Good sea ice close to shore and good ice conditions past Gneiss Point for Cape Roberts work.

Tuesday 18

Prism set erected on flank of Coleman (Neale, Wright, Ashby).

Macpherson began preparation for Cape Roberts work.

Wednesday 19

Prism set on Barnes (Neale, Vardy, Barrett, Ross). Sediment collected from sea ice in New Harbour (Barrett, Ross).

Thursday 20

Drill moved east approximately 200 to 60 m × 40 m ice slab - overthickened 5′5″+.

page 40

Set out ice deflection survey for drill site (Neale, Pyne).

Started moving Drill site building with D4.

Friday 21

Survey: Resection of Mess Hutt at Butter Point Camp. Ice deformation levelling at drill site.

Setting up core saw and other equipment in Science Lab at Butter Point.

Hew motor installed in toboggans for Granite Harbour work.

Saturday 22

Moved survey hut to drill site and set up ice monitoring system. Erected final prism set on Trig Herb for ice monitoring.

Granite Harbour (Macpherson, Wright, Cooper) expedition left but returned after 15 km with broken steering shaft. (Shaft had been rewelded previously).

Shelving installed in Science Lab.

Sunday 23

Survey: Finishing heating system for Rangemaster at drill site.

Moved photography partition in Science Lab.

Picked up Barrett from Scott Base for Granite Harbour.

Monday 24

Granite Harbour expedition left Butter Point (eventually) after starting problems with toboggans.

Drill Site: Deformation survey of sea ice and tested distance measuring equipment.

Built heating parts for tide gauge installation for drill site.

Tuesday 25

Picked up microscope equipment from Scott Base (Pyne, Ashby). Sea ice movement monitoring postponed.

Wednesday 26

Began wiring up Science dry lab. Ward assisting Hoburn.

Started installing tide gauge at Drill Site. Water depth @ 1815 hrs = 196.2 m.

Thursday 27

Completed wiring of Science Labs - now ready for transport and installation at the Drill site.

Granite Harbour return. Sediment traps installed in mid Harbour. Surveying not established because of toboggan unreliability.

Friday 28

Surveying not possible because wheeled vehicles returning to Scott Base in morning. Afternoon snow.

Blasted ice quarry in morning.

Saturday 29

Assisted wiring of CIROS 1 drill site.

Located prism with 2 survey parties @ Herb; and Bernacchi and Coleman with co-ordination with surveyor at drill site. System worked well. First movement monitoring begun today.

page 41

Sunday 30

Set up tide gauge at CIROS I. Sea ice monitoring measurements and levelling at Drill site.

Monday, 1 October

Insulating photography area of Science hut. Loaded 1 hut for move to drill site in afternoon (approximately 1600). Sea ice monitoring measurements.

Tuesday 2 October

Second Science hut moved to drill Bite and began setting up equipment. Sea ice monitoring measurements.

Wednesday 3

Moved fragile science equipment to drill site and continued setting up Science Lab.

Resection surveying at drill site.

Started generator.

Levelling to connect new mark near mud hut with original deformation network at CIROS 1.

Thursday 4

Meeting at Scott Base - to abandon CIROS 1 site and relocate at CIROS 2. Intending to continue monitoring at CIROS 1 for as long as possible.

Friday 5

Began packing equipment and readying Science lab for move to CIROS 2 site.

Located CIROS 2 site and set out CIROS 2 ice deformation network. Strong winds approximately 40 knots from west - network not completed or levelled.

Monitoring at CIROS 1 site.

Retrieved tide gauge from C1 site.

Saturday 6

Completed CIROS 2 deformation network and levelled before site is loaded.

Macpherson and party - preparation for 2nd Cape Roberts trip (Sunday?)

Completed packing of equipment, etc. for move from CIROS 1 to CIROS 2.

Sunday 7

Loaded Science buildings and equipment at C1 site for move to C2. Loading not completed because of high winds at C1 in the evening.

Monday 8

Rangemaster measurements (sea ice movement) at C1.

Tuesday 9

Measured water depth at C2. (209.5 m water level at 1100 hrs).

Science huts moved to C2 and began setting up in late afternoon.

Rangemaster measurements and resection at C1 site.

Wednesday 10

Continued setting up science equipment at C2.

Levelling at C2 and established new level mark.

Thursday 11

Continued with setting up science equipment at C2.

New level marks installed at C2.

Made beacons for New Harbour on shore survey stations.

Rangemaster measurements at C1.

page 42

Friday 12

Reducing and plotting sea ice movement and levelling data.

Continued setting up at C2 and installed tide gauge.

Saturday 13

Started science shift work at C2.

Levelling at C1.

Sunday 14

Coring began, 0.98 m of wash sand retrieved from within HW casing.

Rangemaster measurements at C1 (last set).

Monday 15

API casing - floats retrieved because of drilling problems.

Retrieved prisms from NH1, NH2 sites.

Tuesday 16

Retrieved prisms from C. Bernacchi prism site.

Wednesday 17

API casing - floats are lowered again.

Thursday 18

Core retrieved 0.91 m soft sand-mud, but drilling again in difficulty.

Levelling at C2. Ice thickness at original C1 site now 5′1″.

Friday 19

Drill string stuck - due to bend in casing. No geologist night shift.

Saturday 20

API casing being retrieved.

Prisms from Coleman (last set) retrieved.

Sunday 21

Events K052, K191 brought out to Butter Point. Pick up K041 Snotrac 34 and proceed to Granite Harbour with K042 (Macpherson, Ashby, Vincent). Stop at Spike Cape en route for K052.

Monday 22

Checking K042 Scientific equipment.

Levelling survey at C2 drill site.

Tuesday 23

Herbie - crew stuck at the drill site.

- Granite Harbour party confined to Cape Roberts camp.

Wednesday 24

Thursday 25

Dr. Komura to BP with equipment.

Friday 26

Survey of proposed USGS drill site in Wright Valley (Barrett, Griffith).

Erect drum beacons on trig Herb for New Harbour suveying (Haanen, McLennan, Wright).

Granite Harbour party (K042 and K191) return to Butter Point. Survey of Mackay Glacier Tongue and Cape Roberts sea ice movement survey for CIROS.

Dr. Komura began setting up gas Chromatograph at Drill site.

page 43
Core recovered from Log sheet # Described
6.89 to 8.77 2

Saturday 27

Control points on northern shore of New Harbour surveyed and mast beacons erected (MH2 - Mt Barnes rock, NH6 - C. Bernacchi).

Sunday 28

Ward, Vincent, Macpherson assisting Dunbar (USARP) with the installation of sediment traps in New Harbour.

Core recovered Log sheet # Described
12.14 to 13.06 4
13.20 13.49 5
16.20 16.69 6
20.0 20.77 7

Survey of control points continued.

(NH4 - Hjorth, NH5 - Hjorth-Bern, DVDP 10).

Monday 29

Core recovered Log sheet # Described
27.97 to 28.21 9
28.55 29.17 10

Survey control continued to southern New Harbour (NH1, NH2, F2). CIROS 2 grid levelled.

Tuesday 30

Survey handover Neale - Haanen (reports, computations, data, software, equipment). New Harbour control calculations. Neale depart for SB and NZ.

Wednesday 31

Surveying C. Roberta - Granite Harbour for CIROS associated with K042 until Nov. 4.

Thursday 1 November

Core recovered Log sheet # Described
31.41 to 31.53 11
31.53 32.17 12
32.64 34.42 13
34.47 35.04 14
37.13 38.19 15

Friday 2

Core recovered Log sheet # Described
31.41 to 31.53 16
39.72 40.64 17
40.80 41.86 18
42.05 42.64 19
43.6 45.59 20
46.99 47.28 21
47.43 47.8 22

Saturday 3

Core recovered Log sheet # Described
49.00 to 49.69 23
49.7 51.11 24
page 44

Sunday 4

Core recovered Log sheet # Described
53.15 to 54.40 25
57.75 55.10 26
55.10 55.8 27
56.6 56.83 27
57.6 58.95 28
59.06 61.36 29
61.45 64.54 30
64.59 65.82 31
67.46 68.53 32
68.74 69.83 33
69.83 69.93 34

Monday 5

Core recovered Log sheet # Described
70.06 to 72.26 35
72.18 73.64 36
73.64 76.19 37
76.19 79.02 38
79.05 81.44 39
81.54 83.01 40

Surveying. New drum beacon built. Inner control of southern New Harbour completed (NH1, F2, F3) and beacons left. Position of CIROS 2 co-ordinated. Topographical survey of drill site.

Tuesday 6

Core recovered Log sheet # Described
82.99 to 86.11 41
89.16 92.11 42
92.11 94.40 43
95.60 97.13 44
97.13 98.83 45

Topographical survey of Butter Point Camp. Position of CIROS 1 checked by resection to monitor movement. Ice thickness at CIROS 1 checked.

Wednesday 7

Core recovered Log sheet # Described
98.87 to 101.40 46
101.4 104.43 47
104.4 107.49 48
107.52 110.64 49
110.84 113.25 50
115.9 116.41 51
116.41 116.88 52
117.7 118.5 53
119.47 122.54 54
122.58 125.67 55
125.90 128.87 56
128.87 131.84 57
131.84 134.92 58

Surveying Grid at CIROS 2 relevelled.

page 45

Thursday 8

Core recovered Log sheet # Described
134.91 to 138.0 59
138.12 141.05 60
143.04 144.12 61
146.71 147.4 62
147.72 150.34 63
150.34 153.47 64
153.52 156.61 65
156.67 159.74 66
159.74 162.71 67
169.73 165.76 68
165.76 167.54 69
167.59 168.09 70

Planned helo work on Mt Marston cancelled. Calculations of New Harbour control station co-ordinates. Travelled with Stephenson to SB for mud.

Friday 9

Vehicle caught in blizzard, spent day at SB researching old survey reports - data.

Storm damage at Drill site - coring discontinued.

Saturday 10

Damage assessment at CIROS 2 site.

NQ rods in hole - checked and found tube or overshot could not pass bend at (or near) sea floor. Logging within casing therefore impossible. Decision to begin recovering all drill rod - casing at CIROS 2.

Sunday 11

Begin clean up operation and recovery at CIROS 2.

K042 (Macpherson, Vincent, Cooper) leave for Granite Harbour to set sediment traps. Return 15 November.

Pyne, Stephenson, Griffith to SB to discuss recovery and pull out operation.

Monday 12

Recovery operation CIROS 2 continued.

CIROS 2 site resurveyed for movement (resection).

No significant movement. Levelled deflection grid.

Tuesday 13

Recovery operation continued.

Science Log ends.